OpenStudy (anonymous):

Using complete sentences, explain how you would factor completely x^9 – 64.

OpenStudy (dan815):

x^2-b^2 = (x-b)(x+b)

OpenStudy (usukidoll):

but that's an odd number for the x^9 part

OpenStudy (dan815):

(x^4.5-8)(x^4.5+8)

OpenStudy (usukidoll):

maybe perfect cubes?!

OpenStudy (dan815):

(x^(9/2)-8)((x^(9/2)+8)

OpenStudy (anonymous):

It's a difference of two cubes. Use (a-b)(a^2+ab+b^2).

OpenStudy (usukidoll):

see ^^

OpenStudy (dan815):

u can continue to factor the negative parts

OpenStudy (dan815):

and use complex factors on the positive parts

OpenStudy (usukidoll):

where the heck.... why fractions on exponents the heck?!

OpenStudy (dan815):

HAHAHAHahaa

ganeshie8 (ganeshie8):

it works

OpenStudy (anonymous):

It's still a difference of two cubes, and that's how you're intended to factor it initially. x^9-64 a^3=(x^3)^3 b^3=(8)^3

OpenStudy (usukidoll):

tom and I win!

OpenStudy (anonymous):

Difference of two cubes as in a^3-b^3, of course

OpenStudy (anonymous):

tomhue 64=8^2

OpenStudy (usukidoll):

wait.. cube root of 64 is 4 right? since 4 x 4 = 16 x 4 =64?

OpenStudy (dan815):

64=4*4*4

OpenStudy (anonymous):

Lol sorry, silly mistake. b=4

OpenStudy (cggurumanjunath):